OpenStudy (pastorswife3):

What is the slope of a line passing through points (2,3) and (4,1)?

OpenStudy (anonymous):

-1

OpenStudy (pastorswife3):

Can you show how you got the -1 Sharon504?

OpenStudy (mathstudent55):

For points \((x_1, y_1) \) and \((x_2, y_2) \), the slope of the line that passes through them is \(slope = m = \dfrac{y_2 - y_1}{x_2 - x_1} \)

OpenStudy (mathstudent55):

Subtract the y-coordinates and write the difference in the numerator. Subtract the x-coordinates and write the difference in the denominator. That fraction is the slope. The fraction may be reduced if necessary.

OpenStudy (mathstudent55):

Also, it doesn't matter in which order you subtract the x-coordinates and y-coordinates, but the two subtraction must be in the same order.

OpenStudy (anonymous):

3-1/2-4 2/-2 -1

OpenStudy (mathstudent55):

If you subtract the coordinates of the first point from the coordinates of the second point you get: \(m = \dfrac{1 - 3}{4 - 2} =\dfrac{-2}{2}= -1\) If you subtract the coordinates of the second point from the coordinates of the first point you get: \(m = \dfrac{3 - 1}{2 - 4} =\dfrac{2}{-2}= -1\) As you can see, the point you pick as the first point, and the point you pick as the second point do not alter the result. The slope comes out -1 either way.

OpenStudy (pastorswife3):

Thanks mathstudent55 and sharon504. I now understand!

OpenStudy (mathstudent55):

What you cannot do is subtract one set of coordinates one way and the other set the other way. For example if you subtract the y-coordinates by subtracting the first coordinate from the second, and the when you do the x-coordinates you change the order and subtract the second coordinate from the first coordinate, then the slope will be incorrect. \(\color{red}{This~is~incorrect:} ~~~~~ \color{green}{m = \dfrac{1 - 3}{2 - 4} = \dfrac{-2}{-2} = 1}\) As you can see, by mixing the order of the subtractions, the slope came out 1 instead of -1, and is wrong.